# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v2

@@ B.1 Methodology
B.2 Impact of T on noise-image-latent triangles
Additionally, in Tab. 10 we show how the relations change when sampling and inversion is done for variant $T$.
@@ B.2 Impact of T on noise-image-latent triangles
(a) ADM-32 (CIFAR-10)
(b) ADM-256 (ImageNet)
(c) LDM (CelebA)
B.3 Example histograms
@@
Appendix C Plain surface thresholding
@@
Appendix D Details on calculating inversion error
In Sec. 3.4, we study how the inversion error differs for pixels related to plain and non-plain samples’ regions and investigate it across diffusion steps. In this section, we describe the method for determining the error.
@@ Appendix D Details on calculating inversion error
Further, for the images in the dataset, we obtain binary masks $\mathcal{M}_{p}$ and $\mathcal{M}_{n}$ indicating, respectively, plain ($p$) and non-plain ($n$) pixels in the image, according to the procedure described in Appendix C. To ensure that the level of noise that DM predicts in each step does not bias the results, we divide the absolute errors in each step by $l_{1}$-norm of model outputs, calculated separately for each diffusion step. For plain ($p$) pixels, it can be described as
@@
In Tab. 6, we present the error differences for plain and non-plain areas and how the first $10\%$ of the diffusion steps contribute to the total inversion error. To calculate this error for plain regions in steps $t_{s},\dots,t_{e}$, we sum errors for timesteps from a given interval as
@@
Appendix E Noise-to-image mapping
In Sec. 4.3, we show that the mapping between output images and initial Gaussian noise can be realized through the closest $l_{2}$-distance assignment. In this section, we describe the experiment in detail, indicating this property independence from the number of diffusion steps $T$ and explaining why it does not work in both directions.
E.1 Smallest $l_{2}$ mapping
@@
In our experimental setup, we sample $N=2000$ Gaussian noises and generate $N$ corresponding images using a DDIM sampler. When predicting resulting images for the noises (${\mathbf{x_{T}}}\rightarrow{\mathbf{x_{0}}}$), we iterate over all the $N$ noises, and for each of them, we calculate its pixel distances to all the $N$ generations. For given noise, we select the image to which such $l_{2}$-distance is the smallest. Similarly, we investigate image-to-noise (${\mathbf{x_{0}}}\rightarrow{\mathbf{x_{T}}}$), image-to-latent (${\mathbf{x_{0}}}\rightarrow{\mathbf{\hat{x}_{T}}}$) and latent-to-image (${\mathbf{\hat{x}_{T}}}\rightarrow{\mathbf{x_{0}}}$) mappings.
During the experiment, we calculate the distance between two objects as $l_{2}$ norm of the matrix of differences between them: ${||x-y||}_{2}=\sqrt{\Sigma_{i}^{C}\Sigma_{j}^{H}\Sigma_{j}^{W}(x_{i,j,k}-y_{i,j,k})^{2}}$, where $C\times H\times W$ are the dimensions of either pixel or latent space of diffusion model.
@@
E.2 Impact of the number of diffusion steps on the accuracy of the assignment
[…]
E.3 Asymmetry of Noise-to-Image mapping
Results in Tab. 11(a) indicate that, even though the $l_{2}$-distance is symmetrical, the mapping cannot be done in both directions. The reason behind this is that image and noise assignments are not the same due to the one-directional many-to-one relation, e.g., there might be several noises pointing towards the same closest image.
@@ E.3 Asymmetry of Noise-to-Image mapping
Appendix F Replacing diffusion model predictions during inversion
In Sec. 5 we investigate how Diffusion Model’s prediction approximation error in first inversion steps affect final latent encodings. Here, we present more results for those experiments.
F.1 Impact of replacing diffusion model predictions during inversion on interpolation
In Tab. 12, we show, for DiT model, that replacing the model’s predictions in the first $3$ (6%) inversion steps halves the difference between the quality of interpolations resulting from latents when compared with the noise ones. Additionally, we observe an increase in Recall metric value for middle interpolation generations, suggesting that starting the DDIM inversion from 6% of the image corruption process could significantly increase the usefulness of the latents for interpolation and improve generation diversity.
Inversion predictions SLERPλ=0.5
replaced (%$\%T$) FID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
ø ($0\%$) $19.68$ $0.79$ $0.41$
$1$ ($2\%$) $19.70$ $0.80$ $0.41$
$1\dots 3$ ($6\%$) $18.29$ $0.80$ $0.43$
$1\dots 5$ ($10\%$) $17.51$ $0.81$ $0.43$
$1\dots 10$ ($20\%$) $16.79$ $0.81$ $0.45$
Noise ($100\%$) $16.61$ $0.80$ $0.47$
Table 12: Impact of DDIM Inversion approximation error in first noising steps on the quality of generations resulting from latent interpolations. For the DiT model, we observe that replacing noise approximations with ground truth one in the first $5$ inversion steps leads to higher quality and diversity of interpolated images, as indicated by FID and Recall metrics.
F.2 Impact of replacing diffusion model predictions during inversion on image editing
In Tab. 13, we provide metrics values for Fig. 7. Percentages in each row indicate what fraction of the first inversion steps have been replaced with ground truth DM outputs from the backward diffusion process. Values in Fig. 7 indicate the percentage of improvement in diversity or text alignment for generations in comparison to images sampled from pure latent obtained from standard DDIM Inversion.
Predictions Diversity against $I_{S}$ CLIP Alignment
replaced (%$\%T$) DreamSim $\uparrow$ LPIPS $\uparrow$ SSIM $\downarrow$ DINO $\downarrow$ $P_{S}\downarrow$ $P_{T}\uparrow$
ø ($0\%$) $0.68_{\pm 0.13}$ $0.56_{\pm 0.13}$ $0.26_{\pm 0.15}$ $0.22$ $0.543$ $0.568$
$1$ ($2\%$) $0.68_{\pm 0.13}$ $0.56_{\pm 0.13}$ $0.26_{\pm 0.15}$ $0.22$ $0.471$ $0.667$
$1\dots 3$ ($6\%$) $0.68_{\pm 0.13}$ $0.57_{\pm 0.13}$ $0.26_{\pm 0.15}$ $0.22$ $0.468$ $0.673$
$1\dots 5$ ($10\%$) $0.69_{\pm 0.13}$ $0.57_{\pm 0.13}$ $0.26_{\pm 0.15}$ $0.22$ $0.465$ $0.676$
$1\dots 10$ ($20\%$) $0.69_{\pm 0.13}$ $0.57_{\pm 0.13}$ $0.25_{\pm 0.14}$ $0.20$ $0.462$ $0.681$
Noise $(100\%)$ $0.71_{\pm 0.12}$ $0.59_{\pm 0.12}$ $0.23_{\pm 0.13}$ $0.17$ $0.456$ $0.689$
(a) Diffusion Transformer
Predictions Diversity against $I_{S}$ CLIP Alignment
replaced ($t^{\prime}$) DreamSim $\uparrow$ LPIPS $\uparrow$ SSIM $\downarrow$ DINO $\downarrow$ $P_{S}\downarrow$ $P_{T}\uparrow$
ø ($0\%$) $0.61_{\pm 0.12}$ $0.33_{\pm 0.11}$ $0.41_{\pm 0.14}$ $0.42$ $0.353$ $0.614$
$1$ ($2\%$) $0.61_{\pm 0.12}$ $0.33_{\pm 0.11}$ $0.40_{\pm 0.14}$ $0.41$ $0.352$ $0.615$
$1\dots 3$ ($6\%$) $0.64_{\pm 0.11}$ $0.36_{\pm 0.11}$ $0.38_{\pm 0.13}$ $0.39$ $0.334$ $0.632$
$1\dots 5$ ($10\%$) $0.64_{\pm 0.11}$ $0.36_{\pm 0.11}$ $0.37_{\pm 0.13}$ $0.38$ $0.327$ $0.637$
$1\dots 10$ ($20\%$) $0.65_{\pm 0.10}$ $0.36_{\pm 0.11}$ $0.37_{\pm 0.13}$ $0.37$ $0.313$ $0.645$
Noise ($100\%$) $0.66_{\pm 0.10}$ $0.38_{\pm 0.11}$ $0.34_{\pm 0.14}$ $0.34$ $0.273$ $0.649$
(b) Deepfloyd IF
Table 13: Impact of first DDIM inversion errors on diversity and text-alignment of generations from resulting latent as an input. For IF model (b), replacing the first inversion steps and denoising leads to more diverse generations against the source images $I_{S}$. For DiT (a), we show that such replacing improves the alignment between generation and target prompts, which the generation process is conditioned by, as indicated by the larger CLIP-T value for $P_{T}$ and the smaller one for $P_{S}$.
Appendix G On Noise-Image-Latent relations during diffusion training
@@
G.1 Spatial relations of noise and latents over training time
We conclude our latent localization experiments by showing that our observations are persistent across the diffusion model training process. We generate $N=2048$ images with the final model, using implicit sampling with $T=100$ steps, and invert them to the corresponding latents using checkpoints saved during the training.
@@ G.1 Spatial relations of noise and latents over training time
G.2 Image-to-noise distance mapping over training time
We analyze the image-noise mapping with $l_{2}$-distance from Sec. 4.3 over diffusion model training time. We sample $N=2000$ Gaussian noises and generate images from them using ADM models with $T=100$ diffusion steps, calculating the accuracy of assigning images to corresponding noises (and vice versa) using the smallest $l_{2}$-distance. In Fig. 15, we can observe, for both models, that the distance between noises and their corresponding generations accurately defines the assignment of initial noises given the generated samples (${\mathbf{x_{0}}}\to{\mathbf{x_{T}}}$) from the beginning of the training till the end. At the same time, the accurate reverse assignment (${\mathbf{x_{T}}}\to{\mathbf{x_{0}}}$) can only be observed at the beginning of the training when the trained model is not yet capable of generating properly formed images. Already in the beginning phase of model training, the quality of noise to image mapping rapidly drops and does not change until the end of training.
@@ G.2 Image-to-noise distance mapping over training time
G.3 Image alignment over training time

### v3

@@ F.1 Methodology
(a) ADM-32 (CIFAR-10)
(b) ADM-256 (ImageNet)
(c) LDM (CelebA)
F.2 Impact of T on noise-image-latent triangles
In table 12 we show how the geometrical relations between Gaussian noise, generated images and DDIM latents change when sampling and inversion is done with varying $T$.
@@ F.2 Impact of T on noise-image-latent triangles
F.3 Example histograms
In fig. 11 we present histograms approximating probabiliy density functions for noise ${\mathbf{x_{T}}}$, image ${\mathbf{x_{0}}}$, and DDIM latent ${\mathbf{\hat{x}_{T}}}$ angles.
@@ F.3 Example histograms
Appendix G Plain surface thresholding
@@
Appendix H Calculating inversion error
In section 3.3, we study how the inversion error differs for pixels related to plain and non-plain samples’ regions and investigate it across diffusion steps. In this section, we describe the method for determining the error.
@@ Appendix H Calculating inversion error
Further, for the images in the dataset, we obtain binary masks $\mathcal{M}_{p}$ and $\mathcal{M}_{n}$ indicating, respectively, plain ($p$) and non-plain ($n$) pixels in the image, according to the procedure described in appendix G. To ensure that the level of noise that DM predicts in each step does not bias the results, we divide the absolute errors in each step by $l_{1}$-norm of model outputs, calculated separately for each diffusion step. For plain ($p$) pixels, it can be described as
@@
In table 6, we present the error differences for plain and non-plain areas and how the first $10\%$ of the diffusion steps contribute to the total inversion error. To calculate this error for plain regions in steps $t_{s},\dots,t_{e}$, we sum errors for timesteps from a given interval as
@@
Appendix I Pseudocode for DDIM Inversion with Forward Diffusion
We present in Algorithm 1 the pseudocode for the proposed solution to the decorrelation of latent encodings by replacing the first $t^{\prime}$ inversion steps with a forward diffusion process.
Algorithm 1 Finding decorrelated DDIM latent encoding ${\mathbf{\hat{x}_{T}}}$
1:image $\mathbf{x_{0}}$; diffusion model $\epsilon_{\theta}$; noise schedules $\{\alpha_{t}\}_{t=1}^{T}$, $\{\bar{\alpha}_{t}\}_{t=1}^{T}$; number of inversion steps $T$; forward-replacement timestep $f$
2:decorrelated latent encoding ${\mathbf{\hat{x}_{T}}}$
3:sample $\tilde{\epsilon}\sim\mathcal{N}(0,\mathcal{I})$
4:$\mathbf{\hat{x}_{f}}\leftarrow\sqrt{\bar{\alpha}_{f}}\cdot\mathbf{x_{0}}+\sqrt{1-\bar{\alpha}_{f}}\cdot\tilde{\epsilon}$triangleright forward diffusion
5:for $t\leftarrow f+1,\dots,T$ do
6:     $\mathbf{\hat{x}_{t}}\leftarrow\sqrt{\alpha_{t}}\cdot\mathbf{\hat{x}_{t-1}}+(\sqrt{1-\bar{\alpha}_{t}}-\sqrt{\alpha_{t}})\cdot\epsilon_{\theta}^{(t)}(\mathbf{\hat{x}_{t-1}})$triangleright DDIM Inversion step
7:end for
8:return ${\mathbf{\hat{x}_{T}}}$
Appendix J Details on diversity and alignment of edition
In this section, we provide results for measuring the diversity (against source images $I_{S}$) and alignment to conditioning prompts (source $P_{S}$ and target $P_{T}$) of images generated from Gaussian noise ${\mathbf{x_{T}}}$, DDIM latents ${\mathbf{\hat{x}_{T}}}$, or the latents produced by our method (described in algorithm 1) with modified prompts. We use distance-based metrics (LPIPS [61], DreamSim [11]), similarity metrics (SSIM [57], DINO [5]) to measure the diversity between source (input) images and target (edited) images, as well as the similarity between embeddings produced by the CLIP [44] model to calculate the alignment between the results and prompts. The results in for introduced latent decorrelation procedure are obtained with varying percentages of the first DDIM inversion steps replaced with forward diffusion.
In table 13, we show that by selecting a small fraction of inversion steps to replace with forward diffusion (from $2\%$ up to $6\%$), the resulting latents are more editable.
Inversion steps replaced Diversity against $I_{S}$ CLIP Alignment
by forward diff. (%$\%T$) DreamSim $\uparrow$ LPIPS $\uparrow$ SSIM $\downarrow$ DINO $\downarrow$ $P_{S}\downarrow$ $P_{T}\uparrow$
Noise ${\mathbf{x_{T}}}$ $0.709$ $0.591$ $0.228$ $0.174$ $0.465$ $0.681$
DDIM Latent ${\mathbf{\hat{x}_{T}}}$ $0.677$ $0.564$ $0.261$ $0.223$ $0.480$ $0.662$
$1$ ($2\%$) $0.696$ $0.580$ $0.239$ $0.187$ $0.470$ $0.676$
$1\dots 2$ ($4\%$) $0.701$ $0.587$ $0.227$ $0.179$ $0.469$ $0.677$
$1\dots 3$ ($6\%$) $0.705$ $0.594$ $0.216$ $0.173$ $0.468$ $0.678$
$1\dots 5$ ($10\%$) $0.711$ $0.605$ $0.198$ $0.165$ $0.466$ $0.679$
$1\dots 10$ ($20\%$) $0.722$ $0.628$ $0.166$ $0.153$ $0.465$ $0.681$
$1\dots 50$ ($100\%$) $0.805$ $0.801$ $0.013$ $0.085$ $0.465$ $0.680$
(a) Diffusion Transformer
Inversion steps replaced Diversity against $I_{S}$ CLIP Alignment
by forward diff. (%$\%T$) DreamSim $\uparrow$ LPIPS $\uparrow$ SSIM $\downarrow$ DINO $\downarrow$ $P_{S}\downarrow$ $P_{T}\uparrow$
Noise ($100\%$) $0.665$ $0.380$ $0.339$ $0.344$ $0.273$ $0.649$
DDIM Latent ${\mathbf{\hat{x}_{T}}}$ $0.609$ $0.328$ $0.406$ $0.416$ $0.353$ $0.614$
$1$ ($2\%$) $0.666$ $0.376$ $0.359$ $0.335$ $0.281$ $0.646$
$1\dots 2$ ($4\%$) $0.660$ $0.369$ $0.359$ $0.351$ $0.287$ $0.649$
$1\dots 3$ ($6\%$) $0.661$ $0.370$ $0.353$ $0.349$ $0.285$ $0.650$
$1\dots 5$ ($10\%$) $0.662$ $0.372$ $0.341$ $0.346$ $0.282$ $0.649$
$1\dots 10$ ($20\%$) $0.668$ $0.379$ $0.312$ $0.337$ $0.279$ $0.648$
$1\dots 50$ ($100\%$) $0.733$ $0.521$ $0.004$ $0.272$ $0.274$ $0.649$
(b) Deepfloyd IF
Table 13: Impact of first DDIM inversion errors on diversity and text-alignment of generations from resulting latent as an input. For both DiT (a) and IF (b) models, replacing the first inversion steps and denoising leads to more diverse generations against the source images $I_{S}$. Additionally, we show using forward diffusion in first steps improves the alignment between generation and target prompts, which the generation process is conditioned by, as indicated by the larger CLIP-T value for $P_{T}$ and the smaller one for $P_{S}$.
Appendix K Stochastic editing of real images
In this work, we propose the solution for decorrelating latent encodings resulting from DDIM inversion by replacing its first steps with the forward diffusion. As presented in algorithm 1, the forward diffusion process involves sampling random Gaussian noise $\tilde{\epsilon}\sim\mathcal{N}(0,\mathcal{I})$ and interpolating it with input image. Due to the fact that, when we replace a small fraction of steps ($2-4\%$), the change in image reconstruction error is insignificant, the use of different noises $\tilde{\epsilon}$ (in practice, sampled with different seeds) allows stochastic image editing, i.e. generating different manipulations of the input image, a feature not naturally available with DDIM inversion. In fig. 13, we present examples of editing real images from the ImageNet-R-TI2I dataset (which we annotate using GPT-4o) with the IF model, showing various semantically correct modifications of the same image.
As preserving original image structure during editing is stated as a more difficult task for real images than the one naturally generated by diffusion model, we follow Hertz et al. [14] by, first, denoising latent encodings with source prompt (the one used during inversion), and, after $6\%$ of the steps, using target prompt as conditioning. The examples presented in fig. 13 indicate, that our approach (1) enables stochastic editing of images and (2) enables image manipulations in plain image regions, contrary to editing with naïve DDIM latents.
Figure 13: Replacing first DDIM inversion steps with forward diffusion enables stochastic image editing, resulting in multiple semantically correct manipulations of same input image. Contrary to DDIM Inversion, editing with latents produced by the solution introduced in this work, enables image manipulations in uniform input image areas.
Appendix L Noise-to-image mapping
We showcase an additional study showing the differences that occur between noise and latent encodings, from the perspective of their mapping to the images. Several works investigate interesting properties between the initial random noise and generations that result from the training objective of DDPMs and score-based models. Kadkhodaie et al. [21] show that due to inductive biases of denoising models, different DDPMs trained on similar datasets converge to almost identical solutions. This idea is further explored by Zhang et al. [59], observing that even models with different architectures converge to the same score function and, hence, the same noise-to-imape mapping. Khrulkov and Oseledets [22] show that diffusion models’ encoder map coincides with the optimal transport (OT) map when modeling simple distributions. However, other works [25, 28] contradict this finding.
L.1 Smallest $l_{2}$ mapping
Diffusion models converge to the same mapping between the Gaussian noise (${\mathbf{x_{T}}}$) and the generated images (${\mathbf{x_{0}}}$) independently on the random seed, dataset parts [21], or the model architecture [59]. We further investigate this property from the noise-sample and latent-sample mapping perspective.
@@ Appendix I Pseudocode for DDIM Inversion with Forward Diffusion
In our experimental setup, we start by generating $N=2000$ images ${\mathbf{x_{0}}}$ from Gaussian noise ${\mathbf{x_{T}}}$ with a DDIM sampler and invert the images to latents ${\mathbf{\hat{x}_{T}}}$ with naïve DDIM inversion. Next, we predict resulting images for the starting noise samples (${\mathbf{x_{T}}}\rightarrow{\mathbf{x_{0}}}$) by iterating over all the $N$ noises, and for each of them, we calculate its pixel distances to all the $N$ generations. For given noise, we select the image to which such $l_{2}$-distance is the smallest. Similarly, we investigate image-to-noise (${\mathbf{x_{0}}}\rightarrow{\mathbf{x_{T}}}$), image-to-latent (${\mathbf{x_{0}}}\rightarrow{\mathbf{\hat{x}_{T}}}$) and latent-to-image (${\mathbf{\hat{x}_{T}}}\rightarrow{\mathbf{x_{0}}}$) mappings. We calculate the distance between two objects as $l_{2}$ norm of the matrix of differences between them (with $C\times H\times W$ being the dimensions of either pixel or latent space of diffusion model) as ${||x-y||}_{2}=\sqrt{\Sigma_{i}^{C}\Sigma_{j}^{H}\Sigma_{j}^{W}(x_{i,j,k}-y_{i,j,k})^{2}}$.
[…]
@@
For both noise $x^{T}$ and latents $\hat{x}^{T}$, their assignment to images in both directions can be successfully done when the denoising is performed in the latent space, as shown for DiT and LDM models. We hypothesize that this fact is connected with the KL regularization term that imposes a slight penalty towards a standard normal distribution $\mathcal{N}(0,I)$ on the learned latent during training [45].
L.2 Asymmetry of Noise-to-Image mapping
Results in table 14(a) indicate that, even though the $l_{2}$-distance is symmetrical, the mapping cannot be done in both directions. The reason behind this is that image and noise assignments are not the same due to the one-directional many-to-one relation, e.g., there might be several noises pointing towards the same closest image.
@@
Conclusion.
Those findings, connected with the experiment showing reduced diversity of latents (table 4), suggest that the DDIM latents, unlike noise, cannot be accurately assigned to samples, as the approximation error brings them towards the mean, reducing their diversity and making them closest to most of the images.
Appendix M On Noise-Image-Latent relations during diffusion training
@@ Conclusion.
M.1 Spatial relations of noise and latents over training time
We conclude our latent localization experiments (section 3.2) by showing that our observations are persistent across the diffusion model training process. We generate $N=2048$ images with the final model, using implicit sampling with $T=100$ steps, and invert them to the corresponding latents using checkpoints saved during the training.
@@
M.2 Image-to-noise distance mapping over training time
We analyze the image-noise mapping with $l_{2}$-distance from appendix L over diffusion model training time. We sample $N=2000$ Gaussian noises and generate images from them using ADM models with $T=100$ diffusion steps, calculating the accuracy of assigning images to corresponding noises (and vice versa) using the smallest $l_{2}$-distance. In fig. 17, we can observe, for both models, that the distance between noises and their corresponding generations accurately defines the assignment of initial noises given the generated samples (${\mathbf{x_{0}}}\to{\mathbf{x_{T}}}$) from the beginning of the training till the end. At the same time, the accurate reverse assignment (${\mathbf{x_{T}}}\to{\mathbf{x_{0}}}$) can only be observed at the beginning of the training when the trained model is not yet capable of generating properly formed images. Already in the beginning phase of model training, the quality of noise to image mapping rapidly drops and does not change until the end of training.
@@
M.3 Image alignment over training time
@@
Appendix N Additional results
In this section, we present qualitative results of replacing first DDIM Inversion steps with forward diffusion.
N.1 Failure cases
As described in section 4, while replacing a higher number of inversion steps with forward diffusion enables better editability of latent encodings (through decorrelating them), it increases the image reconstruction error. In fig. 22, we show that replacing $10\%$ or $20\%$ of first steps can introduce significant changes to images.
Figure 22: Replacing first DDIM inversion steps with forward diffusion increases latents editability at cost of a higher reconstruction error. By replacing from $2\%$ up to $4\%$ of steps we obtain reasonable image reconstructions while removing correlations from latents.
N.2 Interpolation
In section 4, we presented that interpolating DDIM latents with SLERP [48] leads to decrease in image quality and diversity when compared to Gaussian noise. In fig. 23, we qualitatively compare our method for removing correlations in latent encodings with naïve DDIM inversion in the task of image interpolation.
Figure 23: Qualitative comparison of images generated from interpolated latents produced with DDIM Inversion and our approach. Contrary to naïve DDIM inversion, proposed solution enables generating high quality objects with pixel-diverse background.